OpenStudy (ratotoseadezhongura):

How would I find a ⋅ b? a = 8i + 6j, b = 4i + 5j I don't think that I can use dot products to find this, and since they are 2D vectors I'm not sure how to use cross products here.

OpenStudy (sshayer):

a.b=(8i+6j).(4i+5j)=8(4)+6(5)=?

OpenStudy (sshayer):

\[a \times b=\left|\begin{matrix}i & j&k \\ 8 & 6&0\\4&5&0\end{matrix}\right|=?\]

OpenStudy (sshayer):

=0i-0j+[8(5)-6(4)}k =16k

OpenStudy (phi):

It looks like a dot product to me. what makes you think it is not ?

OpenStudy (phi):

I treat 8i+6j as completely equivalent to <8,6> i.e. the x component is 8, the "y" component is 6 and <8,6> dot <4,5>= 32+30= 62 (just a scalar number)

OpenStudy (phi):

It looks like a dot product to me. what makes you think it is not ?